OpenStudy (anonymous):

pls help am doing a problem on integration,area between curves y1=x^3, and y2=x which one shouldd be on top?

OpenStudy (anonymous):

which one should be the top boundary?

OpenStudy (anonymous):

I just graphed them and it appears to be that f(x)=x is on top of f(x)=x^3 from obviously 0 to whenever x^3=x. (1)

OpenStudy (anonymous):

oh thanks,and so my points of intersection are 1 and -1?

OpenStudy (anonymous):

Don't forget 0

OpenStudy (anonymous):

how will i use 0 now?

OpenStudy (anonymous):

which boundary should i take 1,0,-1?

OpenStudy (anonymous):

Just keep in mind that they intersect at 0.

OpenStudy (anonymous):

What else does the question/problem say?

OpenStudy (anonymous):

calculate the area bounded by those curves with respect to x axis

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}?\]

OpenStudy (anonymous):

will that be my boundary?

OpenStudy (anonymous):

If you go from -1 to 1, it's just going to be twice as much as whatever the area from 0 to 1 is.

OpenStudy (anonymous):

so can i do it that way,which points do i need to pick?

OpenStudy (anonymous):

problem should have specified a bit more information... but I'd just do 0 to 1.

OpenStudy (anonymous):

Since even if you did do from -1 to 1, you'd have to split up the integral because at some point they switch positions.

OpenStudy (anonymous):

a bit confuse.dont know the boundary to pick now

OpenStudy (anonymous):

anyways i can use 1 and 0 right?

OpenStudy (anonymous):

yeah just do that.

OpenStudy (anonymous):

okay thanks